OpenStudy (anonymous):

A plane flies 450 miles with the wind and 350 miles against the wind in the same length of time. if the speed of the wind is 25mph, what is the speed of the plane in still air? please help!

OpenStudy (anonymous):

450=pt+wt 350=pt-wt where p is the plane's speed and w is the wind, t is time. pt=350+wt, substituting 450=350+2wt then 100=2wt, w=25, so t=2; putting this into an original equation 450=p*2+25*2, 400=2p, p=200 units/unit time

OpenStudy (anonymous):

Thank you very very much!!